# Towards A Universal Graph Structural Encoder

### v1

@@ C.2 Pre-training Metric
    where $Y_{i}^{m}$ is the pre-computed label for the $i$-th node.
Table 8: Test performance of the pre-trained model with different architectures on four pre-training tasks.
Architecture Community Detection Graph Contrastive Learning Shortest Path Distance Motif Counting
Accuracy Accuracy MSE MAE (log)
GatedGCN 0.859 0.831 0.155 2.592
GCN 0.883 0.823 0.141 2.593
GIN 0.892 0.853 0.141 2.611
None+TF 0.828 0.841 0.214 2.603
GCN+TF 0.829 0.841 0.212 2.602
GIN+TF 0.829 0.860 0.193 2.603
None+TF(w/ bias) 0.829 0.863 0.218 2.603
GCN+TF(w/ bias) 0.927 0.857 0.134 2.591
GIN+TF(w/ bias) 0.932 0.862 0.114 2.580
C.3 Biased Attention
The illustration of biased attention is shown in Figure 5
Figure 5: Biased Attention based on random walk matrix.
C.4 Uncertainty-based Loss Weighing
The scale of the loss of different tasks can be different, causing the overall loss to be dominated by a certain task, and ultimately the loss of the other tasks cannot affect the learning process of the network-sharing layers. We use the uncertainty-based loss-weighing method (Kendall et al., 2018) to automatically balance the four pre-training tasks and unify the different scales. Moreover, the uncertainty value reflects the contribution of each task towards the overall pre-training process. A higher uncertainty value indicates a lower contribution (Kendall et al., 2018). Let $\sigma_{\tau}$ and $\mathcal{L}_{\tau}$ represent the task-specific uncertainty value for the task $\tau$. The overall pre-training loss is computed by:
$\mathcal{L}=$ $\frac{1}{\sigma_{\text{SPD}}^{2}}\mathcal{L}_{\text{SPD}}+\frac{1}{\sigma_{\text{MC}}^{2}}\mathcal{L}_{\text{MC}}+\frac{1}{\sigma_{\text{CD}}^{2}}\mathcal{L}_{\text{CD}}+\frac{1}{\sigma_{\text{GCL}}^{2}}\mathcal{L}_{\text{GCL}}$ (12)
C.5 Pre-training Setting
[…]
Appendix D Expressiveness
D.1 Theoretical Proof
For an input graph $G(V,E)$ with node set $V$ and edge set $E$, let $f_{P}:V\rightarrow\mathcal{X}$ and $f_{R}:V\times V\rightarrow\mathcal{X}$ indicate the node-level and edge-level structural encoding, respectively. SEG-WL updates the node labels at the $t$-th iteration by $g_{t}(v)=\text{hash}\left(\left\{\left\{\left(g_{t-1}(u),f_{R}(v,u)\right):u\in V\right\}\right\}\right)$ and $g_{0}(v)=\text{hash}(f_{P}(v))$.
Proposition D.1.
RW($d$)-SEG-WL ($d\geq 3$) is strictly more expressive than WL in testing non-isomorphic graphs.
Proposition D.2.
Proof.
We first introduce Neighbor-SEG-WL, which is the SEG-WL test when $f_{P}$ is an identity encoding and $f_{R}(u,v)$ equals 1 if $(u,v)\in E$ and $2$ otherwise. Previous works have proved the following Proposition (Zhu et al., 2023).
Proposition D.3.
Therefore, Neighbor-SEG-WL is a specific example of SEG-WL test that has equivalent expressiveness to the 1-WL test. We then prove that RW-SEG-WL is strictly more expressive than Neighbor-SEG-WL. Let $d_{\text{neg}}(u,v)$ indicate the edge-level encoding $f_{R}$ in Neighbor-SEG-WL. Note $d_{\text{neg}}(v_{i},v_{j})=2$ if and only if $A_{ij}=0$. Recall that $f_{R}(\cdot,\cdot)$ in RW-SEG-WL satisfies $f_{R}(v_{i},v_{j})=\mathbf{R}_{ij}\in\mathbb{R}^{d}$ with $\mathbf{R}=[\mathbf{I},\mathbf{M},\cdots,\mathbf{M}^{d-1}]$ where $\mathbf{M}=\mathbf{D}^{-1}\mathbf{A}$. Therefore, $f_{R}$ in RW-SEG-WL strictly contains the information of $d_{\text{neg}}$. Therefore, if two non-isomorphic graphs can be distinguished by WL, they can be distinguished by RW-SEG-WL. Proposition D.1 is proved.
To prove Proposition D.2, we provide an example in Figure 6 which shows the Shrikhande graph and the Rook’s $4\times 4$ graph, a pair of strongly regular graphs SRG(16,6,2,2). It is proved that they cannot be distinguished by 3-WL (Arvind et al., 2020). We empirically verified that RW($d$)-SEG-WL with $d>4$ can distinguish these two graphs. ∎
Figure 6: RW($d$)-SEG-WL can distinguish the Shrikhande graph and the Rook’s $4\times 4$ graph when $d>4$ while 3-WL fails
D.2 Synthetic Graph Isomorphism Tests
[…]
Table 9: Results of synthetic graph isomorphism tests
# graphs 21 112 853 11117 15 10 41 180 28 78
# graph pairs 210 6216 363378 61788286 105 45 820 16110 378 3003
number of undistinguishable graph pairs
WL 0 3 17 312 105 45 820 16110 378 3003
SPD-SEG-WL 0 2 12 186 105 45 820 16110 378 3003
RW-SEG-WL($d=8$) 0 0 0 0 0 0 0 0 0 0
D.3 Reconstruction of other PSE types
Table 10: Performance of other PSE reconstruction on $5\%$ MolPCBA dataset. The coefficient of determination $R^{2}$ scores are reported as the metric.
PSE type ElstaticPE LapPE RWSE HKdiagSE CycleSE
GPSE 0.964 0.973 0.984 0.981 0.977
Ours 0.947 0.970 0.987 0.984 0.992
Table 10 demonstrates that the PSE generated by our pre-trained GFSE, followed by a trainable lightweight MLP, is capable of reconstructing various pre-defined PSE types on $5\%$ MolPCBA dataset. We evaluate this using the coefficient of determination $R^{2}$ scores as a metric. Notably, our model performs competitively compared with GPSE, achieving an $R^{2}$ of 0.987 for RWSE and 0.992 for CycleSE, given the fact that GPSE directly adopts PSE reconstruction as its training objective. Instead, our method generalizes well across different PSEs without being directly trained for reconstruction. This suggests that our structural self-supervision tasks are effective and sufficient in capturing important structural information.
Appendix E Evaluation Details
E.1 Integration with Pre-trained Models on Molecules
@@
All models are fine-tuned, trained or tested using five different seeds from 42 to 46, with the results averaged. Additionally, for the results of our generated PSE, we select three different seeds to obtain three GFSE checkpoints. Each GFSE is used to run the downstream task five times with the aforementioned seeds ($42-46$), and all results are averaged.
@@
For the training from scratch setting, we adopt and modify the code base from GPS (Rampášek et al., 2022)  [^1] https://github.com/rampasek/GraphGPS. RWSE and LapPE are of dimension $32$ for molecule benchmark in Table 3 across all the datasets. Given a graph $G(V,E)$, we directly concatenate the PSE and the raw node feature as the new input feature, then send them into the very beginning of a model with randomly initialized parameters, which is as follows:
$X^{\prime}=\operatorname{concat}(X,\text{PSE})$ (13)
$\hat{y}=\operatorname{MLP}[\operatorname{pooling}[\operatorname{GraphModel}(X^{\prime})]],$ (14)
@@
$X^{\prime}=\operatorname{GNN}(X)$ (15)
$\hat{y}=\operatorname{MLP}[\operatorname{pooling}[\operatorname{concact}(X^{\prime},\text{PSE})]],$ (16)
where $X^{\prime}$ denotes the latent node features output from the pre-trained models, PSE denotes the extra structural feature generated by GFSE, the read-out layer consists of pooling and MLPs, where the hyper-parameters follow exactly as  (Xia et al., 2022). We report the performance of the model which achieves the best validation performance during training.
E.2 Integration with Large Language Models
To incorporate the graph structural information into the language model, we use a lightweight MLP to project the $32$-dimensional PSE generated by GFSE into the $4096$-dimensional embedding space of LLaMA2-7B (Touvron et al., 2023). The MLP ensures alignment between the PSE and the language model’s embeddings, facilitating smooth integration of graph information. We then concatenate the textual description of the central node with those of its one-hop neighbors, prepend the projected PSE as a soft token, and append a special graph token at the end of the sequence. This tokenized sequence is fed into the language model, and the hidden embedding corresponding to the special graph token is extracted to represent the central node. We use a contrastive loss between positive node pair $(i,j)\in E$ and negative node pair $(i,j^{\prime})\notin E$ to train the MLP and finetune the language model
$\mathcal{L}=\sum_{(i,j)\in E}\left(d_{ij}^{2}+\max\left(\tau-d_{ij^{\prime}},0\right)^{2}\right),$ (17)
where $d_{ij}=1-\cos(v_{i},v_{j})$ and $v_{i}$ indicates the representation of the node $i$, and $\tau$ is the similarity margin (set as $0.5$ in our experiments).
Appendix F More Experimental Results
F.1 Downstream Evaluation Performance
@@
ogbn-arxiv PubMed MNIST CIFAR10
GateGCN $71.69_{\pm 0.21}$ $76.86_{\pm 0.41}$ $97.34_{\pm 0.14}$ $67.31_{\pm 0.31}$
TF+LapPE $5.86_{\pm 0.00}$ $66.27_{\pm 0.46}$ $96.95_{\pm 0.38}$ $69.01_{\pm 0.61}$
TF+RWSE $5.86_{\pm 0.00}$ $64.43_{\pm 0.37}$ $97.81_{\pm 0.58}$ $70.70_{\pm 0.45}$
TF+GPSE $21.56_{\pm 2.74}$ $65.89_{\pm 0.14}$ $97.78_{\pm 0.32}$ $69.57_{\pm 0.16}$
GPS $70.68_{\pm 0.71}$ $\mathbf{74.26}_{\pm 0.60}$ $98.05_{\pm 0.12}$ $71.49_{\pm 0.35}$
GPS+GFSE $\mathbf{72.30}_{\pm 0.13}$ $74.20_{\pm 0.35}$ $98.15_{\pm 0.46}$ $\mathbf{74.11}_{\pm 0.93}$
[…]
Additional evaluation results on ogbn-Arxiv, PubMed, MNIST and CIFAR10 are given in Table 11, where PubMed and CIFAR10 are out of the pre-training dataset. We observe that most structural encodings surprisingly degrade the performance on the PubMed dataset. The proposed GFSE surpasses other PSE on CIFAR10, demonstrating its generalizability in downstream evaluation.
Table 12: Comparison with GraphPEFT
Cloth Home Sport
Hit@1 MRR Hit@1 MRR Hit@1 MRR
GraphPEFT 76.95 84.71 79.87 86.76 64.61 77.34
   w.o. pre-training 76.74 84.57 79.68 86.63 64.44 77.21
LLaMA + GFSE 76.84 84.68 79.85 86.77 64.79 76.24
Table 12 compares our approach with GraphPEFT. GraphPEFT involves two steps: pre-training a graph encoder on domain-specific graph data and fine-tuning it on evaluation datasets. This process requires a domain-specific encoder for each dataset, increasing the adaptation cost when moving across different domains. In contrast, our model is designed for general-domain usage, offering a more flexible and cost-effective adaptation without requiring domain-specific pre-training. As shown in the table, our model consistently outperforms the version of GraphPEFT that skips the domain-specific pre-training step (w.o. pre-training). This further highlights the robustness and generalizability of our approach, as it avoids the need for costly pre-training on specific domains while still achieving competitive or superior results. Specifically, for the “Sport” dataset, our model demonstrates comparable performance to GraphPEFT with pre-training, further underscoring the adaptability of GFSE in varied contexts.
F.2 Efficiency Evaluation
During the pre-training stage on eight datasets, the average time is around 30 to 40 minutes for each epoch with a single NVIDIA A40 48GB GPU. Total training time is less than two days, which is relatively efficient for a comprehensive multi-dataset pre-training process.
Table 13: Runntimes ($s$) of PSE computation on random synthetic graph with increasing numbers of nodes
@@ F.2 Efficiency Evaluation
We compare the inference efficiency of GFSE with handcrafted positional encodings, such as LapPE and RWSE in Table 13 and Table 14. Specifically, we generate 1,000 synthetic Erdos-Rényi graphs for various graph sizes (100, 300, 500, and 1,000 nodes) and evaluate the time required for pre-computation and inference in GFSE.
Table 14: Runntimes ($s$) on real-world graph dataset
@@
As shown in the table, both LapPE and RWSE exhibit significant increases in computation time as the graph size grows. Pre-computation times required by GFSE inference remain minimal for all graph sizes, underlining the model’s efficiency in this phase. In Table 14, we observe that GFSE demonstrates superior scalability in inference, making it a more efficient option for large-scale graph processing.
Appendix G Discussion
Limitation and Social Impact. While GFSE represents a significant step forward in developing general and expressive foundation models for graph-structured data, there are certain limitations to consider. For example, the effectiveness of GFSE may be influenced by the quality of the pre-training datasets, as biases or under-representation in the data could propagate to the learned representations. Future work could focus on expanding the diversity of pre-training datasets to mitigate such biases and improve the robustness of GFSE across more diverse domains
From a social impact perspective, the development of structural graph encoder like GFSE holds significant promise in advancing various application domains that rely on graph analytics. In fields such as computational biology, social network analysis, and recommendation systems, GFSE could enable more accurate and efficient modeling of complex structured data, leading to improved understanding and decision-making. Furthermore, by reducing the need for extensive task-specific fine-tuning, GFSE could democratize the use of powerful graph learning techniques, making them more accessible to researchers and practitioners with limited computational resources.

### v2

@@ A.4. Pre-training Setting
Appendix B More Experimental Results
B.1. Integration with Pre-trained Models
@@
All models are fine-tuned, trained or tested using five different seeds from 42 to 46, with the results averaged. Additionally, for the results of our generated PSE, we select three different seeds to obtain three GFSE checkpoints. Each GFSE is used to run the downstream task five times with the aforementioned seeds (424642-46), and all results are averaged.
@@
For the training from scratch setting, we adopt and modify the code base from GPS (rampavsek2022recipe)  [^1] https://github.com/rampasek/GraphGPS. RWSE and LapPE are of dimension $32$ for molecule benchmark in Table 4 across all the datasets. Given a graph $G(V,E)$, we directly concatenate the PSE and the raw node feature as the new input feature, then send them into the very beginning of a model with randomly initialized parameters, which is as follows:
$X^{\prime}=\operatorname{concat}(X,\text{PSE});\quad\hat{y}=\operatorname{MLP}[\operatorname{pooling},[\operatorname{GraphModel}(X^{\prime})]]$
@@
(8) $X^{\prime}=\operatorname{GNN}(X);\quad\hat{y}=\operatorname{MLP}[\operatorname{pooling}[\operatorname{concact}(X^{\prime},\text{PSE})]],$
where $X^{\prime}$ denotes the latent node features output from the pre-trained models, PSE denotes the extra structural feature generated by GFSE, the read-out layer consists of pooling and MLPs, where the hyperparameters follow exactly as  (xia2022mole).
B.2. Downstream Evaluation Performance
Table 10. Performance on MolPCBA, ZINC (subset), Peptides-func and Peptides-struct. The best results in each dataset are bolded.
MolPCBA ZINC Peptides-func Peptides-struct
AP $\uparrow$ MAE $\downarrow$ AP $\uparrow$ MAE $\downarrow$
GIN+LapPE $701_{\pm 0.0013} 0.220$$203_{\pm 0.0386}$ $0.5323_{\pm 0.0083}$ $0.2650_{\pm 0.0041}$
GPS+RWSE $2907_{\pm 0.0028} 0.0$$0700_{\pm 0.0040}$ $0.6603_{\pm 0.0085}$ $0.2739_{\pm 0.0063}$
GPS+GPSE $2911_{\pm 0.0036} 0.0$$0648_{\pm 0.0030}$ $0.6688_{\pm 0.0151}$ $\mathbf{0.2464}_{\pm 0.0025}$
@@ B.2. Downstream Evaluation Performance
GFSE Imp.($\%$) $32.60$ $76.43$ $2.78$ $42.47$
From Table 10, we observe that the optimal selection of structural encoding typically varies across different datasets and base models. For example, RWSE tends to be more beneficial than LapPE for small molecular graph learning (e.g., MolPCBA and ZINC), whereas the opposite is generally true on Peptides. Notably, the performance gains are most pronounced when integrating PSE with Transformer, demonstrating its critical role in compensating for the absence of inherent structural sensitivity in Transformers. The last row shows the average improvement ($\%$) brought by GFSE on base models. The consistent improvements across different settings underscore the robustness and generalizability of GFSE, especially in contexts where traditional PSE fails to deliver.
B.3. Efficiency Evaluation
During the pre-training on eight datasets, the average time is around 30 to 40 minutes for each epoch with a single NVIDIA A40 48GB GPU. Total training time is less than two days, which is relatively efficient for a comprehensive multi-dataset pre-training process.
Table 11. Runntimes ($s$) of PSE computation on random synthetic graph with increasing numbers of nodes
@@ B.3. Efficiency Evaluation
We compare the inference efficiency of GFSE with handcrafted positional encodings, such as LapPE and RWSE in Table 11 and Table 12. Specifically, we generate 1,000 synthetic Erdos-Rényi graphs for various graph sizes (100, 300, 500, and 1,000 nodes) and evaluate the time required for pre-computation and inference.
Table 12. Runntimes ($s$) on real-world graph dataset
@@
As shown in the Table 11, both LapPE and RWSE exhibit significant increases in computation time as the graph size grows. Pre-computation times required by GFSE inference remain minimal for all graph sizes, underlining the model’s efficiency in this phase. In Table 12, we observe that GFSE demonstrates superior scalability in inference, making it a more efficient option for large-scale graph processing.